Two fair dice are rolled. Find the probability that the sum of the two numbers is not greater than 5.
Answer by stanbon(48568)

Two fair dice are rolled. Find the probability that the sum of the two numbers is not greater than 5.
-----------
Random variable values are 2,3,4
2-combinations: 1,1
3-combinations: 1,2; 2,1
4-combinations: 1,3; 3,1; 2,2
5-combinations: 1,4; 4,1; 2,3; 3,2
--------------
There are 36 sum combinations
--------
Probability (sum not greater than 5)=10/36=5/18
Cheers,
Stan H.